Mountain View Company produces hiking boots. The direct labor standard for each pair of boots is 1 hour at a cost of $16.00 per direct labor hour. During the month of May, Mountain View Company used 3,000 direct labor hours to produce 2,250 pairs of boots. Total direct labor cost for May was $42,000. What is the labor rate variance for May? OA. $6,000 unfavorable OB. $6,000 favorable O C. $4,500 unfavorable O D. $4,500 favorable
Answer #1

Solution:

Actual rate of labor = $42,000 / 3000 = $14 per hour

Labor rate variance = (SR - AR) * AH = ($16 - $14) * 3000 = $6,000 Favorable

Hence option B is correct.